### v3

# WeChat-YATT: A Scalable, Simple, Efficient, and Production Ready Training Library

Abstract
Reinforcement Learning from Human Feedback (RLHF) has emerged as a prominent paradigm for training large language models and multimodal systems. Despite the notable advances enabled by existing RLHF training frameworks, significant challenges remain to scale to complex multimodal workflows and adapt to dynamic workloads. In particular, current systems often encounter limitations related to controller scalability when managing large models, as well as inefficiencies in orchestrating intricate RLHF pipelines, especially in scenarios that require dynamic sampling and resource allocation. In this paper, we introduce WeChat-YATT (Yet Another Transformer Trainer in WeChat), a simple, scalable, and balanced RLHF training framework specifically designed to address these challenges. WeChat-YATT features a parallel controller programming model that enables flexible and efficient orchestration of complex RLHF workflows, effectively mitigating bottlenecks associated with centralized controller architectures and facilitating scalability in large-scale data scenarios. In addition, we propose a dynamic placement schema that adaptively partitions computational resources and schedules workloads, thereby significantly reducing hardware idle time and improving GPU utilization under variable training conditions. We evaluate WeChat-YATT across diverse experimental scenarios, demonstrating its substantial throughput improvements over state-of-the-art RLHF training frameworks. Furthermore, WeChat-YATT has been successfully deployed to train models that support WeChat product features for a large-scale user base, underscoring its effectiveness and robustness in real-world applications. We have made WeChat-YATT publicly available at https://www.github.com/tencent/WeChat-YATT.
Keywords Distributed Systems  $\cdot$ Deep Learning  $\cdot$ Generative Reward Modeling  $\cdot$ Reinforcement Learning from Human Feedback
1 Introduction
Transformer-based large models and their multimodal extensions have achieved remarkable progress, demonstrating state-of-the-art performance across a wide range of natural language processing, understanding, and generation tasks [1, 2, 3, 4, 5, 6, 7, 8, 9]. These models, characterized by their massive parameter counts and ability to process extremely long context sequences, have become the foundation for many advanced AI applications. However, the unprecedented scale of LLMs has introduced significant challenges for efficient training, and traditional data-parallel training strategies are often insufficient to handle the memory and computational demands of such large models. To address these challenges, the research community has developed a variety of hybrid parallelism techniques [10, 11, 12, 13, 14]. These approaches enable the training of models that would otherwise exceed the memory and compute capabilities of individual devices, by partitioning both the model and the data across multiple GPUs or nodes.
The integration of Reinforcement Learning from Human Feedback (RLHF) has significantly transformed the training paradigms of large language models [15, 16, 17]. By embedding human preferences into the optimization process, RLHF requires the coordination of multiple models—such as reward models, policy models, and reference models—within a unified training pipeline [18, 19, 20]. This multi-model architecture substantially amplifies the number of parameters and the volume of intermediate activations that must be handled, thereby straining the capabilities of current distributed training infrastructures. The scalability challenges are further exacerbated in multimodal RLHF scenarios. The incorporation of diverse data modalities (e.g., text, images, audio) not only increases the heterogeneity and volume of training data but also intensifies the demands on inter-node communication and memory bandwidth. As a result, the system faces pronounced communication overheads and memory bottlenecks, hindering its ability to scale efficiently to larger datasets and model sizes. Addressing these bottlenecks is critical for facilitating the practical deployment of RLHF in increasingly complex and data-rich environments.
Sample efficiency is a critical factor in the RLHF training process, as it directly influences convergence rates and overall model performance [21]. While flexible sampling strategies have been shown to accelerate convergence, most existing RLHF training frameworks lack robust support for efficient and adaptive sampling mechanisms. Although hybrid parallelism theoretically enables substantial scalability, practical implementations often encounter significant communication overhead as the degree of parallelism increases. Specifically, as the number of models and the length of context sequences expand, the costs associated with synchronizing parameters and activations across distributed devices can escalate rapidly, thereby diminishing the anticipated benefits of parallelization. To address these challenges, recent work such as Shen et al. [22] has conceptualized the RLHF training workflow as a control flow problem, employing techniques such as multi-cluster, multi-model partial colocated, and memory swapping to alleviate peak GPU memory consumption. Similarly, other systems [23, 24] have explored alternative strategies, with Sheng et al. [24] further enhancing throughput by integrating both single- and multi-controller architectures alongside statically optimized placement schemes. These advancements underscore the ongoing need for innovative approaches to improve sampling efficiency and resource utilization in large-scale RLHF training.
  1. 1.
    Scalability bottlenecks in large-scale multimodal scenarios: The management of extensive multimodal data (e.g., images, videos) can result in a single controller becoming a communication or memory bottleneck, thereby constraining system throughput or even causing operational failures.
  2. 2.
    Inefficiency under dynamic sampling and generative reward modeling: Workflows that require frequent dynamic sampling or generative reward computation are susceptible to significant overhead from model swapping and long-tail effects (i.e., straggler tasks), which can substantially reduce hardware utilization and training efficiency.
To address these challenges, we introduce WeChat-YATT, a novel RLHF training framework designed to enhance scalability, flexibility, and efficiency in large-scale, multi-model training environments. The key contributions of this work are as follows:
  1. 1.
    Parallel Controller Programming Model: We propose a parallel controller programming model that facilitates the construction of complex RLHF workflows. This model supports efficient distributed execution and flexible control flow, effectively mitigating the bottlenecks associated with centralized controller architectures.
  2. 2.
    Dynamic Scaling Placement Schema: We present a dynamic scaling placement schema that enables efficient generative rewarding and dynamic sampling. By adaptively partitioning resources and supporting fine-grained scheduling, our approach minimizes device idle time (“bubbles”) and optimizes hardware utilization, even under highly variable workloads.
The proposed WeChat-YATT advances the state-of-the-art in RLHF training by effectively addressing the critical bottlenecks of controller scalability and resource placement. We hope that WeChat-YATT will serve as a robust foundation for both academic research and industrial applications, fostering further innovation and collaboration within the community.
2 Background
2.1 Distributed Training of Deep Models
Data Parallelism involves splitting the training data across multiple devices (such as GPUs), where each device holds a copy of the model and processes a different mini-batch of data. Gradients are averaged (or summed) across devices to update the model parameters synchronously [25, 26]. Tensor parallelism splits individual layers or tensors of the model across multiple devices. Each device is responsible for computing a portion of the operations within a layer, enabling the training of very large models that cannot fit into a single device’s memory [10]. Pipeline parallelism divides the model into sequential stages, with each stage placed on a different device. Mini-batches are divided into micro-batches and passed through the pipeline, allowing different devices to work on different parts of the input simultaneously, improving hardware utilization [11, 27, 28].
ZeRO is an optimization technique that partitions model states (such as optimizer states, gradients, and parameters) across devices to minimize memory redundancy. This enables the training of much larger models by reducing the memory consumption per device [29, 30].
Context Parallelism is a parallelization strategy where input context is distributed across multiple devices. This allows for efficient processing of long sequences or large batches by splitting the workload. Techniques like ring attention are examples of context parallelism, enabling scalable and memory-efficient attention computation across devices [2, 14, 31, 32].
Expert Parallelism is used in Mixture-of-Experts (MoE) models, where different subsets of the model (experts) are distributed across devices. During training and inference, only a subset of experts is activated for each input, allowing efficient scaling and utilization of resources [13, 33].
2.2 RLHF Workflow
RLHF aligns the linguistic space of LLMs with human preferences, using a set of human-ranked candidates of given prompts [18, 34]. An typical reinforcement learning trainer consists of multiple models, e.g., an actor, a critic, a reference policy, and one or multiple reward models. The critic and reward models are usually models tuned in the human preference dataset, with the language modeling head replaced by a numerical output head (Bradley-Terry reward model) [34, 35, 36, 37]. Bradley-Terry reward models can be replaced with generative models, which obtain reward scores through token prediction; furthermore, leveraging chain-of-thought reasoning, the accuracy and final performance of the actor model can be further improved [38, 39].
Most RLHF workflows largely follow this 4-stage workflow [18, 19, 40, 16, 15, 41, 20](Sometimes, the first three stages are referred to as ’rollout’ (Figure 1(a)):
  1. 1.
    Stage 1(Generation): The actor produces responses from a batch of prompts using auto-regressive generation.
  2. 2.
    Stage 2(Rewarding): The reward model computes rewards for generated responses.
  3. 3.
    Stage 3(Preparation): Using generated responses and rewards, the critic computes their values, the policy and reference policy compute their reference log probabilities.
  4. 4.
    Stage 4(Training): The actor and the critic are trained on the batch of data produced by previous stages and the loss function.
(a) Regular RLHF workflow
(b) RLHF workflow with Dynamic Sampling
Figure 1: RLHF workflow
In engineering practice, the generation step is typically implemented using a specialized language model inference engine (e.g., vLLM [42], sglang [43]) [23, 22, 24].
To simplify implementation, most systems use a single controller to manage data transfer, intermediate results, and workflow orchestration through a central point [23, 24]. However, this centralized approach can sometimes become a bottleneck, especially when the trainer itself has high memory requirements.
2.3 RLHF Placement
Workflow is at the core of RLHF training, and model placement is a crucial component of the RLHF workflow. Currently, there are at least two main approaches:
  • Co-exist: Different models are assigned to separate groups of devices, allowing them to run in parallel as long as there are no data dependencies between them.
  • Co-locate: Multiple models are placed on the same set of GPUs, sharing GPU memory. To avoid out-of-memory (OOM) errors, which can easily occur if large models are run simultaneously, these colocated models are executed one after another in a time-sharing fashion.
Although Sheng et al. [24] proposed methods for calculating placement based on workload, both the open source implementations of OpenRLHF [23] and VeRL [24] adopt a co-locate all models strategy. The partial colocated approach introduces overhead from swapping among multiple models. For example, swapping a 32B model could take nearly a minute, and updating weights may take tens of seconds. In typical GRPO training, this overhead is not a major issue, since the primary bottlenecks in RLHF training are the generation and the forward/backward passes of modeling. Compared to the tens of minutes spent on rollout and training, model swapping is not the system bottleneck.
Some works [44, 45] have adopted a coexisting approach, which effectively avoids the overhead of model swapping. However, to mask device idle time during workflow state transitions, these methods introduce varying degrees of asynchronous overlap between the rollout and training stages. This can lead to suboptimal model convergence; for example, Fu et al. [44] introduces staleness-aware training to address this issue. In our experience, many algorithm engineers find that this trade-off adds extra cognitive burdens.
3 Design of WeChat-YATT
3.1 Overview
Figure 2: Overview of partial colocated WeChat-YATT and colocated WeCaht-YATT
Figure 2 presents a high-level overview of the proposed WeChat-YATT framework. WeChat-YATT supports two distinct resource placement architectures: partial colocated and full colocated.
In the partial colocated architecture, the Sampler and the GenRM (Generative Reward Model) components are deployed independently and interact asynchronously. The allocation of computational resources to these two components directly influences throughput during the rollout phase, a placement method that will be examined in detail in Section 3.3. Within this architecture, the actor model leverages all available machine resources across the cluster, while parallel controllers correspond to the ranks of model-parallel worker heads associated with the Actor. The training loop is initiated by wake-up remote procedure calls (RPCs), activating both the Sampler and GenRM. Subsequently, parallel controllers retrieve data from the distributed GDataset, thereby mitigating excessive memory consumption on any single machine. Notably, both the Sampler and GenRM operate asynchronously: upon completing each microbatch, the Sampler immediately transmits the generated data to the GenRM without impeding ongoing generation processes. Once the rollout phase concludes, the Actor instructs other components to enter a sleep state, offloads their memory, and subsequently loads the actor model to initiate the training phase.
In contrast, the full colocated architecture adopts a more integrated approach, wherein all components share the resources of all machines. Under this configuration, the GenRM must await the Sampler’s completion of the entire sample generation process. Following this, the Sampler offloads its parameters to disk, and the GenRM is loaded into memory to initiate the reward computation process. Upon completing reward computation, the Actor is loaded into memory, while the GenRM is offloaded. This workflow is inherently synchronous, with transitions between components occurring only after the preceding process is fully completed. Consequently, in scenarios characterized by frequent long-tail effects, the full colocated architecture may not offer optimal performance.
Overall, WeChat-YATT’s flexible resource placement strategies enable efficient orchestration of RLHF workflows, accommodating diverse operational requirements and workload characteristics.
3.2 Parallel Controllers
While hybrid controllers offer notable advantages in terms of programming convenience, they may lack the necessary flexibility or introduce system bottlenecks under specific conditions. Two principal challenges inherent to both single and hybrid controller paradigms merit attention.
Resource Bottlenecks in Data-Intensive Workflows. A significant limitation arises when large-scale features—such as extensive collections of images or videos, particularly in image generation tasks [5, 7, 46]—must be transferred within the control flow. In such cases, the memory capacity and RPC network bandwidth of a single controller can quickly become limiting factors. Additionally, the execution of complex procedures within the controller may result in CPU bottlenecks. For instance, in specialized algorithmic scenarios, even with a controller provisioned with 2048 GB of memory, practical constraints such as model and optimizer swap buffers typically reduce available memory by half. Conducting a rollout of 1024 samples, each comprising 32 images at 2K resolution, would already consume approximately 768 GB of memory. Further operations, including data copying, increased video frame counts, or communication buffer overhead, can precipitate out-of-memory (OOM) errors. Employing shallow copies can mitigate memory consumption by preventing redundant tensor creation during repeated sampling. Nevertheless, if the sampling algorithm incorporates advanced data augmentation strategies, memory usage may escalate once again.
Limited Granularity in State Transitions. From a systems architecture perspective, the hybrid controller essentially functions as a single entity managing proxy objects for GPU resource pools associated with multiple roles (e.g., policy, reward, reference, critic). This design restricts the controller to stage-wise transitions, thereby impeding the realization of local state changes. The inability to perform localized state transitions complicates support for specialized sampling processes, such as dynamic sampling or reward-augmented generation [40, 47].
To address these challenges, we introduce a parallel controller architecture. As illustrated in Figure 3, reinforcement learning (RL) tasks are partitioned using the conventional SPMD (Single Program Multiple Data) paradigm, wherein multiple controllers independently manage distinct data subsets. Statistically, as batch size increases, the distribution of workload among controllers becomes increasingly balanced. This effect is underpinned by the law of large numbers, which ensures that, with sufficiently large sample sizes, the data allocated to each controller converges toward its expected value, thereby promoting equitable load distribution.
In the parallel controller architecture, each controller is responsible for managing a subset of system resources, with individual resources potentially managed by either a single controller or shared among multiple controllers. This design enables the concurrent existence of multiple operational stages within the system, thereby facilitating fine-grained state transitions that can be tailored to the requirements of specific reinforcement learning algorithms.
(a) Hybrid Controller
(b) Parallel Controllers
Figure 3: Each controller in parallel controllers is only responsible for part of workload and resources, avoids overwhelming of controller bottleneck resources
Coordinating multiple reinforcement learning roles using several controllers is often considered a complex undertaking. In our proposed framework, however, we allocate distinct controllers to manage different RL roles. This approach mirrors the hybrid controller paradigm [24], wherein each actor internally employs a multi-controller structure to enhance computational efficiency [48, 42].
The parallel controller architecture can be interpreted as a pragmatic evolution of the hybrid controller, designed to provide increased controller resources and enable more flexible resource scheduling. For programming efficiency, our experience suggests that, the adoption of a multi-controller architecture does not pose substantial programming difficulties. In our implementation, we further decompose the top-level controller and employ collective communication mechanisms to facilitate coordination among controllers. Within each worker cluster managed by a controller, we preserve a control pattern analogous to that of the hybrid controller, thereby maintaining operational consistency while improving scalability and flexibility.
3.3 Dynamic Placement
The generative rewarding approach has demonstrated considerable effectiveness in recent studies [38, 49]. In contrast to discriminative LLM-based verifiers, which do not leverage the generative capabilities of pretrained LLMs, generative rewarding utilizes standard next-token prediction to assess solution correctness. Specifically, this method represents correctness through the LLM’s probability distribution over tokens, rather than by predicting a separate numerical score. This design preserves the generative abilities of the language model, as the verification decision is formulated as an additional token prediction. Furthermore, it enables several inherent advantages of LLMs, including unified training for both solution generation and verification, support for chain-of-thought reasoning, and the ability to perform inference-time computation [49].
To implement this, we adopt a partial colocation strategy, which, as previously discussed, offers an efficient solution without introducing additional algorithmic trade-offs. In our framework, a causal text generation inference engine replaces the traditional regression-based rewarding model. Upon completion of the generation phase, the policy model is offloaded to CPU memory, and the generative reward model is loaded onto GPU memory. This reward model then generates reward scores through a combination of text generation and regular expression matching.
Empirically, the partial colocation approach has proven highly effective, resulting in minimal hardware idle time. When compared to the relatively processes of text generation, reward computation, and model training, the overhead associated with swapping models in and out of GPU memory is negligible. For example, swapping a 32B model typically requires only 20 seconds for SGLang, which constitutes a minor fraction of the total computation time.
However, dynamic sampling scenarios introduce additional complexity. As described in Yu et al. [40], prompts with accuracy scores of 1 or 0 (i.e., batches with uniformly high or low rewards) are filtered out, triggering resampling. This process necessitates additional rounds of generation and reward computation, as illustrated in Figure 1(b).
  1. 1.
    Frequent Swapping Overhead. As training progresses and model performance improves, the acceptance rate of generated samples tends to decrease. This results in more frequent resampling and, consequently, more frequent model swaps between the policy and reward models. Under these conditions, the previously negligible model swapping overhead can accumulate, potentially becoming a bottleneck in the training pipeline and limiting overall system efficiency.
  2. 2.
    Long-tail Accmulation. During the generation phase, the occurrence of long-tail outputs is inevitable. These outputs can reduce overall GPU cluster utilization. In partial colocation settings, where frequent transitions between multiple stages are required, the long-tail effect may be further exacerbated. Specifically, the uneven completion times of different tasks can lead to increased resource fragmentation and idle GPU time, thereby posing additional challenges to maintaining high cluster efficiency.
To address the challenges associated with model swapping overhead and the exacerbated long-tail effect, we develope a dynamic placement strategy that integrates dynamical resource allocation in the training process under patial colocated WeChat-YATT paradigm.
In stages 1 and 2, the simultaneous residency of both the policy generation model and the rewarding model in GPU memory eliminates the need for model swapping, thereby avoiding the memory copy and graph capturing overhead typically incurred during dynamic sampling. Furthermore, by maintaining the policy model in memory following the reward computation, subsequent rounds of generation can be initiated immediately. This finer-grained control effectively reduces idle periods ("bubbles") during the long-tail phase of the system, thereby enhancing overall hardware utilization and system efficiency.
From an engineering standpoint, dynamically partitioning the GPU cluster according to workload is technically feasible. However, this approach does not fully account for the fact that system load can fluctuate due to algorithmic factors [16]. During reinforcement learning, for example, the average response length of models on the training set tends to increase as the model learns to solve more complex reasoning tasks, often requiring extended "thinking time." This trend is consistently observed throughout the training process, as models progressively acquire the capability to address increasingly sophisticated reasoning challenges by leveraging longer test-time computations. Such computations may involve generating responses that range from hundreds to thousands of reasoning tokens, thereby enabling the model to explore and refine its reasoning processes in greater depth. Our simulation experiments have frequently exhibited phenomena consistent with these observations.
When the workload is dynamic, static estimation methods such as those proposed in Sheng et al. [24] become inadequate for determining optimal resource placement. It is inherently difficult to predict how the output length of the model will evolve over the course of training. Moreover, as the model’s reasoning responses become longer, it becomes increasingly challenging to anticipate the corresponding generation length required for reward computation. This unpredictability further complicates the design of efficient placement strategies in dynamic training environments.
To mitigate the aforementioned challenges, we propose a dynamic placement strategy. At the onset of training, initial resource allocation between the policy model and the generative rewarding model is determined using simple heuristic methods. As training progresses, hardware utilization is continuously monitored, enabling the gradual reduction of resources assigned to roles exhibiting low utilization and the reallocation of these resources to roles with higher demand. This adaptive approach ultimately leads to a balanced workload distribution across training roles and maximizes hardware utilization.
Empirical observations indicate that generation time exhibits a unimodal relationship with respect to the resources allocated to a single component (e.g., the Sampler). To optimize resource placement, we employ the Ternary Search algorithm, as outlined in Algorithm 1. The algorithm is executed at regular intervals to ensure that the placement strategy remains optimal throughout the training process.
Input: Real generation time $f(x)$ with $x$ GPUs for sampler, number of GPUs $n$
Output: optimal $x^{*}$
$r=n,l=1$
while rl>0r-l>0 do
 $m_{1}\leftarrow l+\frac{r-l}{3},m_{2}\leftarrow r-\frac{r-l}{3}$
 if $f(m_{1})>f(m_{2})$ then
    $l\leftarrow m_{1}$
 else
    $r\leftarrow m_{2}$
$x^{*}\leftarrow l$
return $x^{*}$
Algorithm 1 Ternary Search for Dynamic Placement
4 Implementation
We implement WeChat-YATT using Python and PyTorch [50]. Our system combines vLLM [42] and SGlang [43] for generation serving with Megatron-Core [10, 11, 12, 13] as the training backend.
4.1 Multi-processing
No software is perfect; every piece of software inevitably has some flaws or limitations. Some programs may create conflicting global objects, or rely on global state that makes multi-tenancy impossible. These kinds of issue can lead to unexpected behavior, or difficulties when integrating different components. This can cause subtle or hard-to-detect issues when running multiple models in a single process. For example, some distributed training libraries use a global communication process group, making it impossible to assign different hybrid parallel plans to different models for optimal throughput. Due to the internal scheduling of enterprises, it is unrealistic to expect the open-source community to deliver bug fixes or undertake large-scale refactoring within a specific timeframe. This reality necessitates that enterprises plan ahead and develop their own solutions, ensuring that their projects can proceed according to their own schedules without being hindered by external dependencies.
We distribute all modules, whether for generation, training, or rewarding, across different processes, and enable them to collaborate via Remote Procedure Call. At the same time, we strive to minimize interference with the orchestration mechanisms used by these modules. By running each component in a separate process, we can avoid issues where multiple instances of certain components interfere with each other, which could otherwise impact throughput or introduce subtle bugs. Furthermore, by respecting the original design of each component and minimizing our interference with their orchestration, we allow the software to operate along well-tested code paths as much as possible. This approach not only improves stability and reliability, but also makes it easier to diagnose and resolve issues when they arise.
4.2 Placement and RPC
To keep the placement scheme as simple and predictable as possible (e.g., forming communication groups according to the GPU switch topology) and to minimize interference with component orchestration, we did not use Ray [51] for placement and RPC. Instead, we launch tasks via WeChat’s internal job scheduling system. Nevertheless, our code is designed to be loosely coupled, and we believe it can interoperate well with systems such as MPI [52] and SLURM [53]. We do not doubt that Ray could also accomplish similar tasks.
To ensure RPC delivery and exactly-once semantics, we designed a simple and generic RPC mechanism: each RPC request is assigned a unique ID, and the result is cached on the server side until the client successfully retrieves it. The client then sends a request to clean up the cached RPC result. Since deep learning training systems typically only consider complete success (with all other cases treated as complete failure), error handling is greatly simplified. If the RPC returns an unexpected or undesired result, the controller simply terminates all processes.
Since it is impossible to completely avoid rare and untested code paths in components that may cause the system to hang, or low-probability scenarios within the internal cluster that may result in extremely slow CPU execution, we monitor the training progress to ensure it meets expectations. If the training progress falls below the expected threshold, the job is terminated, resources are reallocated, and the job is restarted.
4.3 Idle Resources and Checkpointing
Most WeChat services exhibit strong temporal fluctuations in resource usage, resulting in a large amount of idle resources during off-peak hours. We leverage these idle resources for non-urgent model training tasks.
To minimize training progress loss due to interruptions, we employ asynchronous checkpointing to improve checkpoint efficiency and increase checkpoint frequency. Additionally, when online services detect increased load and request resources from the training cluster, we attempt to save an on-demand checkpoint. If the checkpoint cannot be completed within the specified time, we abandon the current progress and release resources to prioritize online service needs.
To accommodate elastic resource scaling, we utilize distributed checkpointing and design the dataloader consumption state such that checkpoints can be reused across GPU clusters of varying sizes.
4.4 Workload Balancing
As noted in Wang et al. [54] and Jiang et al. [55], workload imbalance can create significant bubbles during training, becoming a major bottleneck. When sequence lengths are long, the main training bottleneck is attention computation, whose complexity is $s^{2}$ for a sequence of length $s$, while the cost of packing is much lower. For example, if a sequence is formed by concatenating two sequences of length $\frac{s}{2}$, the total computation is $\frac{s^{2}}{2}$, leading to severe load imbalance and device underutilization. This issue is even more pronounced with post-training data, which often varies greatly in length.
Compared to complex combinatorial optimization approaches, we find a much simpler solution for workload balancing: instead of sequence packing, we simply sort data by simulated workload ($s^{2}$), greatly simplifying the process and even improving computational efficiency.
Most of the time, the length distribution of the data is normal. This means that, in practice, the workload is also normally distributed, which makes it easy for us to construct global batches with very little padding. In the worst case, if the data lengths are randomly distributed, the computational waste is still minimal. Suppose the maximum length is $s$, and the local reordering batch buffer size is $B$, then the useless computation within a single batch does not exceed $1-\left(\frac{B-1}{B}\right)^{2}$. For example, when $s=1M,B=16$, at most 8.6% of the computation is wasted.
Naively sorting data by workload(length) may introduce distribution bias. To address this, we first bucket data according to the global batch size, then shuffle the buckets to ensure data is randomly distributed. Our experiments show that this method has almost no impact on model accuracy.
4.5 Distributed Attention
Currently, most implementations achieve context parallelism using ring attention for both causal and full attention. To support more complex attention masks (e.g., Gemma-3 [56]), inspired by Llama-3 [2], we implement distributed attention based on CCL all-gather, where we first all-gather the key (K) and value (V) tensors, and then compute attention output for the local query (Q) tensor chunk.
To reduce the memory footprint of all-gathered key-value pairs, inspired by open source community contributors, we process only a subset of attention heads at a time and overlap KV communication with attention computation. This makes it feasible to train sequences up to 1 million tokens in length.
4.6 Train Data Storage
Multimodal data is often extremely large, both in quantity and size [57, 56]. While algorithm developers typically use JSONL files for flexible storage of text data, images and videos cannot be easily stored in this format.
Storing massive numbers of images directly in a distributed file system can easily exceed file number quota, putting pressure on the distributed file system name nodes. To address this, we adopt private key-value storage solutions (e.g. FeatureKV and UnionDB), on top of our private service discovery and distributed file system (Wechat File System) [58], to serve training data. By leveraging more efficient storage engine data structures, we achieve more efficient and convenient storage and access for training data.
5 Evaluation
We conduct our experiments on a cluster of NVIDIA GPUs, where each node is equipped with eight NVIDIA H20-96GB GPUs interconnected via NVLink. The inter-node communication leverages RDMA with a bandwidth of 200 Gbps. The software stack comprises CUDA 12.4, PyTorch 2.6, Megatron-core 0.12.2, and SGLang 0.4.6.post5. While the evaluation presented in this paper utilizes a relatively small number of GPUs to validate WeChat-YATT, it is noteworthy that large-scale GPU clusters—comprising up to thousands of GPUs—are deployed in WeChat’s internal production environment. However, due to compliance constraints, further details regarding the production deployment cannot be disclosed.
5.1 Generative Reward Model as Critic in Training
We begin by evaluating WeChat-YATT in conjunction with a generative reward model on a mathematical reasoning task. Specifically, we employ the Qwen 2.5 model series [15] as the backbone models and utilize the GSM8K dataset [59] for training.
We conduct experiments with the following methods:
  • VeRL[24]: The framework is an efficient reinforcement learning platform designed to support scalable, distributed training and evaluation of RL algorithms across diverse environments proposed be seed team.
  • Full Colocated WeChat-YATT: This framework is implemented over WeChat-YATT in a full colocate manner. All components in the framework are colocated across clusters sharing all resources.
  • Partial Colocated WeChat-YATT: In this framework, a partial colocation strategy is employed to facilitate dynamic sampling.The actor is colocated with other components, while the sampler—responsible for actor rollouts—operates independently alongside the generative reward model.
The metric of the evalutaion is execution time in unit of seconds to reflect the overall throughput of different framework.
The comparative evaluation of the Generative Reward Model as Critic in Training is conducted in two parts. First, we assess the throughput of the generative reward model under a full colocated WeChat-YATT deployment compared with sota framework VeRL. Subsequently, we investigate the impact of dynamic sampling in specific scenarios, where frequent swapping introduces additional overhead and exacerbates long-tail latency.
5.1.1 Co-locate Generative Reward Model
In the full colocated WeChat-YATT setup, we conduct experiments on the generative reward model compared with VeRL under two distinct scenarios. The first scenario features a balanced parameter size between the generative reward model and the actor model, while the second scenario considers a highly unbalanced configuration. As VeRL currently supports only external calls to the generative reward model, we allocate computational resources in proportion to the parameter size ratio between the generative reward and actor models.
For the first comparation scenario setup, we select the VeRL Generative Reward demo. In this configurationof VeRL, GenRM-CI-Test-1.5B serves as the external reward model, utilizing 8 GPUs for inference. SGlang is employed as the inference engine, with a data parallelism size (dp-size) of 8. An additional 8 GPUs are allocated to deploy the actor model, Qwen2.5-3B-Instruct, with a tensor parallelism size (tp-size) of 2. For the training backend, we adopt the state-of-the-art Megatron-LM framework. To ensure a fair comparison, the WeChat-YATT setup mirrors the VeRL configuration, employing a full colocated deployment with 16 GPUs.
(a) Training Time of Full Colocated WeChat-YATT and VeRL
(b) Time consumption breakdown
Figure 4: Training time and time consumption breakdown with GenRM-CI-Test-1.5B as gen-rm, and Qwen2.5-3B-Instruct as actor
As illustrated in Figure 4, WeChat-YATT achieves a 59.2% reduction in average training time compared to VeRL. This improvement is primarily attributed to WeChat-YATT’s full colocated strategy, which enables the utilization of all available GPUs throughout the training process. Specifically, WeChat-YATT offloads the parameters of the inference engine during training and reloads them as needed, thereby minimizing idle periods for both the generative reward model and the actor model. To further investigate the performance advantage of WeChat-YATT in this scenario, we analyze four key stages of a PPO step. As shown in Figure 4(b), each stage benefits from the partial colocated strategy, as all 16 GPUs are actively engaged due to the offload-onload mechanism, resulting in consistently faster execution across all stages.
(a) Training Time of Full Colocated WeChat-YATT and VeRL
(b) Time consumption breakdown
Figure 5: Training time and time consumption breakdown with Qwen2.5-Math-72B-Instruct as gen-rm, and Qwen2.5-Math-1.5B as actor
We further evaluate performance in a scenario involving a large generative reward model, Qwen2.5-Math-72B-Instruct, paired with a smaller actor model, Qwen2.5-Math-1.5B. This configuration corresponds to the generative reward model demonstration provided in our open-source codebase. For the VeRL setup, 24 GPUs are allocated for inference, while an additional 8 GPUs are dedicated to actor model rollout generation and training. In contrast, WeChat-YATT employ a full colocated configuration utilizing 32 GPUs, leveraging an offload/onload strategy. Megatron-LM is used as the training backend in both cases.
As depicted in Figure 5, WeChat-YATT achieves a 20.0% reduction in average training time across PPO training steps, as detailed in Figure 5(a). In this extreme scenario, the primary bottleneck is the generative reward model evaluation, as shown in Figure 5(b). While WeChat-YATT optimizes the actor rollout generation and training phases, this comes at the cost of additional overhead associated with offloading and onloading the large generative reward model. Notably, even though WeChat-YATT utilizes 32 GPUs for the generative reward model—compared to 24 GPUs in the VeRL setup—it exhibits a slightly longer reward evaluation time. This is attributable to the dominance of onload-offload overhead in scenarios characterized by low sequence length and small rollout batch size, where the additional communication and memory management costs outweigh the benefits of increased GPU parallelism. Nevertheless, WeChat-YATT accelerates the remaining stages of the training pipeline, resulting in a lower overall training time compared to VeRL.
Table 1: Rollout Time Usage under Different Sample Rejected Rates
10% 20% 30% 40%
Total (s) Avg. (s) Total (s) Avg. (s) Total (s) Avg. (s) Total (s) Avg. (s)
Full Colocate 1987.1 39.7 2350.6 47.0 2838.4 56.8 3457.3 69.1
Partial Colocate 1509.1 30.1 1968.5 39.4 2315.9 46.3 2684.4 53.7
5.1.2 Dynamic Sampling
To assess the performance of the partial colocated WeChat-YATT framework in a dynamic sampling context, we conduct experiments in which varying sample rejection rates are applied during the rollout phase. Consistent with previous experimental settings, GenRM-CI-Test-1.5B is employed as the generative reward model, while Qwen2.5-3B-Instruct served as the actor model. The sample rejection rates are varied from 10% to 40%, and the framework need to re-sample rejected rollouts until all samples are accepted. The primary metric of interest is the time consumption observed during the rollout phase under these different rejection rates.
As shown in Table LABEL:tab:reject_rates, the rollout time for both frameworks increases as the sample rejection rate rises. However, the partial colocated WeChat-YATT consistently demonstrates lower rollout time across all rejection rates when compared to the full colocated approach. Specifically, the partial colocated framework achieves a reduction in rollout time usage ranging from 16.2% to 24.1% relative to the full colocated configuration.
This performance improvement can be attributed to the asynchronous interaction between the actor rollouts and the generative reward model in the partial colocated design. In this setup, each completed microbatch of actor rollouts is immediately dispatched to the standalone generative reward model in an asynchronous manner. This process effectively mitigates the long-tail latency that often constrains throughput in the full colocated WeChat-YATT framework. Furthermore, the frequent context switching between actor rollouts and the generative model in the full colocated setup introduces additional overhead, which is substantially reduced in the partial colocated architecture.
5.2 Dynamic Placement
Figure 6: Dynamic placement algorithm generation time compared with Weighted-Based placement
We adopt the same experimental setup as described in subsection 5.1.2. In this set of experiments, we vary the generation sequence length to examine how generation time changes as the sequence length increases. For comparison, we employ the Weighted-Based placement strategy, in which computational resources are allocated in proportion to the model weights of the generative reward model and the actor model.
The experiments results are presented in Figure 6. The reported generation time metric represents the average across all PPO steps. As expected, our results demonstrate that WeChat-YATT consistently achieves lower generation times than the weighted-based resource placement strategy across different generation sequence lengths.
The superior performance of WeChat-YATT can be attributed to its ability to identify the optimal resource placement strategy during training via the ternary search algorithm. Furthermore, by partially colocating both the policy and generative reward models in GPU memory, WeChat-YATT effectively eliminates swapping overhead. This approach also substantially mitigates the long-tail effect, further enhancing system efficiency.
6 Conclusion
In this work, we present WeChat-YATT, a simple, scalable, and balanced RLHF training framework that addresses critical bottlenecks in large-scale, multi-model reinforcement learning workflows. By introducing a parallel controller programming model and a dynamic scaling placement schema, WeChat-YATT effectively mitigates memory and communication bottlenecks, and adapts to dynamic workloads with improved hardware utilization. Our design emphasizes practicality and flexibility, aiming to reduce engineering complexity while supporting efficient orchestration of complex RLHF pipelines.
WeChat-YATT has already produced various models that deployed in WeChat’s production environment, supporting real-world business applications that serve the entire user base. This practical deployment demonstrates the framework’s robustness and scalability. Although challenges remain in optimizing resource allocation and accommodating increasingly diverse model architectures, we believe WeChat-YATT offers a strong foundation for advancing research and deploying large-scale, human-aligned models in practice.